OpenStudy (horsegal244):

Charlie's Computer Company charges $0.65 per pound to ship computers. Part A: Write an equation to determine the total cost, c, to ship p pounds of computers. Use your equation to determine the cost of shipping 2 pounds of computers. Part B: If the company reduces the cost to ship computers by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of computers with the reduced cost.

OpenStudy (horsegal244):

@eta

OpenStudy (horsegal244):

@Hero

OpenStudy (horsegal244):

@Michele_Laino

OpenStudy (rainbow_rocks03):

plz I need help on a question

OpenStudy (horsegal244):

Well I need Help too

OpenStudy (horsegal244):

@kjones0331

OpenStudy (z4k4r1y4):

thin about this... if it costs $0.65 to ship 1 pound of weight how much does it cost to ship 2 pounds?

OpenStudy (horsegal244):

$1.30

OpenStudy (horsegal244):

@Z4K4R1Y4

OpenStudy (michele_laino):

it is simple, we have: c=0.65*p

OpenStudy (michele_laino):

since we have to pay $0.65 for one pound

OpenStudy (z4k4r1y4):

so you multiplied the amount of pounds (p) to 0.65 and you got the cost (c) so you have c=0.65p

OpenStudy (horsegal244):

Ok now what

OpenStudy (z4k4r1y4):

so for part (A) you have 'p' = 2. and for part (B) it's the same equation but instead of 0.65 you have 0.05.

OpenStudy (michele_laino):

part B: the new cost per pound is now: 0.65-0.05=0.60

OpenStudy (z4k4r1y4):

sorry i mad a mistake Michele is right.

OpenStudy (z4k4r1y4):

*made

OpenStudy (michele_laino):

so the new equation, is: c= 0.60*p

OpenStudy (horsegal244):

ok now what

OpenStudy (z4k4r1y4):

Finished, thats all you need to do for part B.

OpenStudy (horsegal244):

ok

OpenStudy (horsegal244):

can u help with 2 more

OpenStudy (horsegal244):

@Z4K4R1Y4

OpenStudy (z4k4r1y4):

I can try.